OpenStudy (anonymous):

A car rental agency charges $16.00 per day plus $.15 per mile. Jim's bill for 4 days was $79.00. How many miles did he drive?

OpenStudy (mathstudent55):

Cost = (fee per day) * (number of days) + (cost per mile) * (number of miles) 79 = 16 * 4 + 0.15 * m 79 = 16*4 + 0.15m 0.15m + 64 = 79 0.15m = 15 m = 15/0.15 m = 100 Answer: 100 miles

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

Eight times what number is greater than 72?